the ratio of men to women working in a company is 2 to 3. if there are 81 women working for the company what is the total number of employees
Answer by stanbon(75887)

the ratio of men to women working in a company is 2 to 3. if there are 81 women working for the company what is the total number of employees
----
2:3 is the same as 2x:3x where 3x is # of women.
---
3x = 81
x = 27
-----------
# of men = 2x = 2*27 = 54
----
Total # of employees = 54+81 = 135
=====================================
Cheers,
Stan H.
==============